OpenStudy (anonymous):

Find the LCM of r^2-81 and r^2+14r+45

OpenStudy (anonymous):

r^2-81=(r+9)(r-9) r^2+14r+45=r^2+5r+9r+45 =(r+5)(r+9) so the lcm is (r+5)(r+9)(r-9)